OpenStudy (anonymous):

give the parametrization for the circle with radius 5 and center (-2,4) that for 0

OpenStudy (anonymous):

A circle has the equation (in rectangular coordinates) \[(x-h)^2+(y-k)^2=r^2\] where the center is \((h,k)\) and the radius is \(r\). You're given that the circle has radius 5 and center (-2,4), which means it has the equation (again, in rectangular) \[(x+2)^2+(y-4)^2=25\] We want to be able to describe every point \((x,y)\) using a different parameter \(t\), so that every x- and y-coordinate of each point on the circle can be described in terms of \(t\). In other words, we're looking for functions \(x(t)\) and \(y(t)\) that, for a given value of \(t\), give a point that lies on the circle.

OpenStudy (anonymous):

For circles, it'll help to recall some trigonometry. Here's the Pythagorean identity: \[\cos^2t+\sin^2t=1\] Now this doesn't match up perfectly with what we want to express, so we have to make the appropriate adjustments. Multiply both sides by 25, then you have \[25\cos^2t+25\sin^2t=25\] Still not quite there. Let's rewrite a bit more: \[(5\cos t)^2+(5\sin t)^2=25\] Now notice that if we replace \(5\cos t\) with \(x+2\) and \(5\sin t\) with \(y-4\), we get our circle: \[(\color{red}{5\cos t})^2+(\color{blue}{5\sin t})^2=25~~\iff~~(\color{red}{x+2})^2+(\color{blue}{y-4})^2=25\]

OpenStudy (anonymous):

So basically we've substituted \[\begin{cases} x+2=5\cos t\\ y-4=5\sin t \end{cases}~~\iff~~\begin{cases} x=5\cos t-2\\ y=5\sin t+4 \end{cases}\] This is your parameterization. However, it's not *quite* what you're looking for. You're given that the \(t\) is taken from \(0\le t\le2\pi\). But when we let \(t=0\), we have \[\begin{cases} x=5\cos 0-2\\ y=5\sin 0+4 \end{cases}~~\iff~~\begin{cases}x=3\\y=4\end{cases}\] So for \(t=0\), we get the point \((3,4)\), but we want to get \((-7,4)\). Notice also that if we take \(t=\dfrac{\pi}{2}\), we have \[\begin{cases} x=5\cos \dfrac{\pi}{2}-2\\\\ y=5\sin \dfrac{\pi}{2}+4 \end{cases}~~\iff~~\begin{cases}x=-2\\y=9\end{cases}\] So we have \((-2,9)\) for \(t=\dfrac{\pi}{2}\).

OpenStudy (anonymous):